# Athena: Enhancing Multimodal Reasoning with Data-efficient Process Reward Models

### v1

@@ 5 Conclusion
Appendix A More Experiments and Details
A.1 Choices of Completers
Here we verify the effect of combining different completers on the accuracy of estimated labels. Results are listed in Table 6. From Table 6, it is noted that the consistency filter for two completers usually improves the accuracy of estimated labels. For example, combining two weak or strong completers improves accuracy by almost two points in general. Furthermore, the accuracy of estimated labels is improved by combining one weak and one strong completer, achieving the most competitive accuracy, e.g., 95.2% with weak completer $\phi_{w}^{2}$ and strong completer $\phi_{s}^{1}$. This shows that the proposed consistency filter strategy is effective across all combinations of different completers and achieves the best performance when we combine a weak completer and a strong completer.
@@ A.2 Benchmarks
MathVista [36] Testmini 1000 Accuracy
MathVision [54] Full 3040 Accuracy
MathVerse [68] Vision-Only 788 Accuracy
DynaMath [70] Full 5050 Accuracy
WeMath [44] Testmini 1740 Score (Strict)
LogicVista [60] Full 448 Accuracy
MMMU [65] Validation 900 Accuracy
GSM8K [9] Test 1319 Accuracy
MATH [18] Test 5000 Accuracy
VisualProcessBench [57] - 2866 F1-score
A.3 Reward Choice
In this paper, we consider three reward choices of Athena-PRM for aggregating step scores into a final score [69], including PRM-minium, PRM-last, and PRM-product. For all reasoning steps $\{a_{i}\}_{i=1}^{K}$, its corresponding reward for each step is $\{r_{i}\}_{i=1}^{K}$, the reward $r$ for each solution with different scoring methods are calculated as follows: 1) PRM-minium uses the minium reward across all steps, i.e., $r=\min\{r_{i}\}_{i=1}^{K}$; 2) PRM-last uses the reward of last step, i.e., $r=r_{K}$; 3) PRM-product uses the reward product of all steps, i.e. $r=\prod_{i=1}^{K}r_{i}$.
@@ A.3 Reward Choice
Table 8: Best-of-8 performance comparison on the MathVista [36] with Qwen2.5-VL-7B [3] with different reward choices: last step reward, minium reward across all steps and product of all rewards. Default settings are marked in gray.
@@
A.4 Label Distribution
We present the distribution of process labels in Table 9. It is shown that label imbalance exists in PRM800K and Athena-300K. We try different up-sample rates (e.g. ×$\times 2$) for samples with error steps and find that slightly up-sampling negative data is beneficial for training PRMs, as shown in Table 10. We find that increasing the up-sample rate will not improve performance. We set the up-sample rate to 2 for other experiments.
@@ A.4 Label Distribution
Athena-PRM ×$\times 2$ 75.2
Athena-PRM ×$\times 5$ 75.2
Athena-PRM ×$\times 10$ 74.9
Appendix B Limitations and Social Impacts
Limitations.  Athena-PRM achieves impressive performance across various scenarios and benchmarks. However, Athena-PRM is specifically designed to enhance the reasoning capabilities of LLMs and MLLMs. Its application may be limited to solving problems with clear answers, such as multi-step reasoning and math problems. Effectively assigning rewards for intermediate steps in open-ended text generation tasks remains an open problem, and we leave this for future research.
[…]
Social Impacts.   The paper introduces an effective method for generating high-quality, process-labeled data for reasoning problems. Our approach enhances the efficiency and effectiveness of training PRMs. We hope that our research will have a broader impact on improving the reasoning abilities of foundation models. Since our method focuses on general improvements in the reasoning capabilities of foundation models, it does not directly produce negative societal impacts.
Appendix C Case Study
@@
Hint: Please answer the question and provide the correct option letter, e.g., A, B, C, D, at the end. Question: When the ant <image1> walks from home <image2> along the arrows 3,3,3,1$\rightarrow$ 3$\uparrow$ 3$\rightarrow$ 3$\uparrow$ 1→ 3 , ↑ 3 , → 3 , ↑ 1, he gets to the ladybird <image3>. Which animal does the ant <image1> get to when he walks from home <image2> along the following arrows: 2,2,3,3,2,2$\rightarrow$ 2$\downarrow$ 2$\rightarrow$ 3$\uparrow$ 3$\rightarrow$ 2$\uparrow$ 2→ 2 , ↓ 2 , → 3 , ↑ 3 , → 2 , ↑ 2? <image6> <image7> Choices:(A) A (B) B (C) C (D) D (E) E (Answer: A)
@@
Figure 3: A case study from VisualProcessBench [57].
@@ 3. Angle properties:
Figure 4: A case study from VisualProcessBench [57].
@@
[…]
Figure 5: A case study from MathVista [36] with Qwen2.5-VL-72B [3].
@@
[…]
Figure 6: A case study from MathVista [36] with Qwen2.5-VL-72B [3].

### v2

@@ 4 Conclusion
Appendix A The Use of Large Language Models (LLMs)
We only use LLMs to polish our drafts, and LLMs do not evolve research ideation.
Appendix B Related Work
Test-time scaling.  While scaling data and model parameters during training has been extensively studied (Kaplan et al., 2020; Hoffmann et al., 2022), scaling computation at test-time has garnered interest more recently (Snell et al., 2025; Setlur et al., 2025). TTS enables language models to leverage additional computational resources when faced with challenging questions. Our study primarily focuses on parallel scaling (Zhang et al., 2025a), wherein language models generate multiple outputs simultaneously and aggregate them into a final answer. This aggregation can involve selecting the most common outcome (Wang et al., 2023) or using reward models (Snell et al., 2025) to rank solutions, selecting the one with the highest reward, as demonstrated in our Best-of-N settings.
Process reward models.  Reward models are crucial in reinforcement learning (Guo et al., 2025; Shao et al., 2024; Wang et al., 2024b; Ouyang et al., 2022; Bai et al., 2022; Dong et al., 2024) and test-time scaling (Setlur et al., 2025; Lightman et al., 2024; Wang et al., 2024b; Cobbe et al., 2021; Snell et al., 2025). Outcome reward models (Cobbe et al., 2021) assign a scalar value to question-response pairs, whereas process reward models evaluate each step, typically achieving superior performance and generalization (Lightman et al., 2024; Luo et al., 2024; Setlur et al., 2025; Li & Li, 2025). While better performance and generalization are achieved, collecting data to train PRMs is challenging. PRM800K (Lightman et al., 2024) is the first open-source process supervision dataset annotated by humans, but scaling it is challenging due to the time and skill required for annotating reasoning steps. Math-Shepherd (Wang et al., 2024b) proposes an automated method for estimating intermediate step correctness using Monte Carlo estimation, though it generates some incorrect labels and demands extensive computational resources. Our work explores strategies to reduce computational costs while improving process label accuracy.
Appendix C More Experiments and Details
C.1 Choices of Completers
Here we verify the effect of combining different completers on the accuracy of estimated labels. Results are listed in Table 6. From Table 6, it is noted that the consistency filter for two completers usually improves the accuracy of estimated labels. For example, combining two weak or strong completers improves accuracy by almost two points in general. Furthermore, the accuracy of estimated labels is improved by combining one weak and one strong completer, achieving the most competitive accuracy, e.g., 95.2% with weak completer $\phi_{w}^{2}$ and strong completer $\phi_{s}^{1}$. This shows that the proposed consistency filter strategy is effective across all combinations of different completers and achieves the best performance when we combine a weak completer and a strong completer. We also explore use more completers to further improve the qualify of process labels. However, we did not observe any significant improvement, despite the increased computational cost.
@@ C.2 Benchmarks
MathVista (Lu et al., 2024b) Testmini 1000 Accuracy
MathVision (Wang et al., 2024a) Full 3040 Accuracy
MathVerse (Zhang et al., 2024) Vision-Only 788 Accuracy
DynaMath (Zou et al., 2025) Full 5050 Accuracy
WeMath (Qiao et al., 2025) Testmini 1740 Score (Strict)
LogicVista (Xiao et al., 2024) Full 448 Accuracy
MMMU (Yue et al., 2024) Validation 900 Accuracy
GSM8K (Cobbe et al., 2021) Test 1319 Accuracy
MATH Hendrycks et al. (2021) Test 5000 Accuracy
VisualProcessBench (Wang et al., 2025) - 2866 F1-score
C.3 Reward Choice
In this paper, we consider three reward choices of Athena-PRM for aggregating step scores into a final score (Zhang et al., 2025b), including PRM-minium, PRM-last, and PRM-product. For all reasoning steps $\{a_{i}\}_{i=1}^{K}$, its corresponding reward for each step is $\{r_{i}\}_{i=1}^{K}$, the reward $r$ for each solution with different scoring methods are calculated as follows: 1) PRM-minium uses the minium reward across all steps, i.e., $r=\min\{r_{i}\}_{i=1}^{K}$; 2) PRM-last uses the reward of last step, i.e., $r=r_{K}$; 3) PRM-product uses the reward product of all steps, i.e. $r=\prod_{i=1}^{K}r_{i}$.
@@ C.3 Reward Choice
Table 8: Best-of-8 performance comparison on the MathVista (Lu et al., 2024b) with Qwen2.5-VL-7B (Bai et al., 2025) with different reward choices: last step reward, minium reward across all steps and product of all rewards. Default settings are marked in gray.
@@
C.4 Label Distribution
We present the distribution of process labels in Table 9. It is shown that label imbalance exists in PRM800K and Athena-300K. We try different up-sample rates (e.g. $\times 2$) for samples with error steps and find that slightly up-sampling negative data is beneficial for training PRMs, as shown in Table 10. We find that increasing the up-sample rate will not improve performance. We set the up-sample rate to 2 for other experiments.
@@ C.4 Label Distribution
Athena-PRM $\times 2$ 75.2
Athena-PRM $\times 5$ 75.2
Athena-PRM $\times 10$ 74.9
C.5 Effectiveness of Athena-PRM for Reward Ranked Fine-tuning
Besides selecting responses with Athena-PRM, we have tried to directly use all responses to fine-tune. We do not use Athena-ORM to select responses because all data we use in reward-ranked fine-tuning has ground-truth answers. Before we use all data to fine-tune, we conduct a quick test to verify the effectiveness of response ranking with Athena-PRM. We used a small-scale dataset ( 30k queries) to verify the effectiveness of response ranking with Athena-PRM. Results are shown in Table 11.
Table 11: Results of reward ranked fine-tuning (RAFT) with different data settings. We report accuracy on the MathVista with Qwen2.5-VL-7B.
Method Data Athena-PRM Accuracy
Qwen2.5-VL-7B - - 68.1
+RAFT 30k 68.7
+RAFT 30k 69.8
+RAFT (ours) 150k 71.4
Appendix D Case Study
@@
Hint: Please answer the question and provide the correct option letter, e.g., A, B, C, D, at the end. Question: When the ant <image1> walks from home <image2> along the arrows $\rightarrow 3,\uparrow 3,\rightarrow 3,\uparrow 1$, he gets to the ladybird <image3>. Which animal does the ant <image1> get to when he walks from home <image2> along the following arrows: $\rightarrow 2,\downarrow 2,\rightarrow 3,\uparrow 3,\rightarrow 2,\uparrow 2$? <image6> <image7> Choices:(A) A (B) B (C) C (D) D (E) E (Answer: A)
@@
Figure 3: A case study from VisualProcessBench (Wang et al., 2025).
@@ 3. Angle properties:
Figure 4: A case study from VisualProcessBench (Wang et al., 2025).
@@
Answer: $\boxed{60}$ 0.04
Figure 5: A case study from MathVista (Lu et al., 2024b) with Qwen2.5-VL-72B (Bai et al., 2025).
@@
Answer: $\boxed{6}$ 0.99
Figure 6: A case study from MathVista (Lu et al., 2024b) with Qwen2.5-VL-72B (Bai et al., 2025).